# Token Distillation: Attention-Aware Input Embeddings for New Tokens

### v2

@@ 5.2 In-Depth Analysis
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Original tokenization 93.4 88.1 99.2 91.1 97.8 65.2 90.9 94.0 90.0 Random (Hewitt, 2021) 0.0 0.0 0.0 0.0 0.0 0.0 0.2 0.0 0.0 Mean (Gee et al., 2022) 2.2 4.2 9.1 8.0 11.3 4.2 7.4 7.4 6.7 NTP (Lampinen et al., 2018) 34.0 54.3 70.6 49.7 71.6 29.0 53.9 70.0 54.1 ZeTT (Minixhofer et al., 2024) 15.3 25.4 28.4 – – – – – – Token Distillation 46.3 68.0 86.3 61.8 81.3 31.6 57.5 84.3 64.6 Token Distillation + NTP 48.3 68.0 85.3 60.2 79.7 38.2 63.4 80.9 65.5 Token Distillation + NTP 46.9 71.0 85.7 62.0 80.1 37.2 66.4 83.7 66.6
@@
In Table 4, we report LLM-as-a-Judge results on the correctness of generated definitions. This tests the limits of our method, as the new tokens are now more complicated, such as “Software Development” or even common phrases such as “spill the beans”. In this setting, even when using the original tokenization, the model sometimes is not able to generate correct definitions. In turn, the gap of Token Distillation to the original tokenization is also larger in this experiment. However, Token Distillation outperforms all other baselines by a very large margin. ZeTT (Minixhofer et al., 2024) in particular struggles in the multi-word token setup, likely because it is out-of-distribution from the hyper-network pretraining.
@@
We find that the subtoken mean initialization as starting point yields better results than random initialization (see “Random + Token Distillation” in Table 7). If a pretrained embedding prediction hyper-network is available, we can instead use these predicted embeddings as the starting point. We run this experiment using ZeTT; this combination further improves performance compared to using the subtoken mean (see Table 7). We also compare refining the hypernetwork embeddings with NTP. This greatly improves over NTP but yields only minimal gains over using the hypernetwork embeddings themselves without further NTP.
@@
Our main proposed method Token Distillation computes the distillation objective based on a mean-squared error (MSE) of last layer hidden states. A natural question is whether we can also use the “traditional” knowledge distillation objective (Hinton et al., 2015) based on the Kullback-Leibler divergence (KL-Div) (Kullback & Leibler, 1951). We analyze this and report the results in Table 6. We additionally include computing the MSE between logits instead of hidden states. For the KL-Div based results, we include different variations of combining the objective with a NTP-based objective. In general, the choice of distillation objective does not matter much. This is hardly surprising, as the logits are merely a linear projection (potentially with an additional normalization layer) of the hidden state; the log-probabilities used for KL-Div are then in turn merely an additional log-normalization via log-softmax. Combining the KL-based results with NTP-based objectives follows similar trends as their combinations with hidden state MSE-based objectives. We opt for the MSE on hidden states for two reasons: (1) In terms of implementation convenience, the objectives with a projection to logits over the vocabulary require a masking step for the “student” model, as the original “teacher” model does not have the new tokens in its vocabulary. (2) Choosing hidden states allows us to explore other layers than just the last one with potentially large computational speedups – we explore this next.
@@
In our main experiments, we apply the Token Distillation objective on the hidden states after the last layer. However, the last hidden state might not be the most optimal, as it might be overspecialized for next-token prediction (Rogers et al., 2020). In Figure 2, we plot the average performance on our biomedical benchmark datasets using Llama3-8B-Instruct while varying the target layer for extracting hidden states for our Token Distillation objective. Indeed we do see a slight downward trend as we approach the last layer – choosing a different layer than the last one can further boost our already strong results. We repeat the analysis with Llama3.1-8B-Instruct in Section A.2 with matching results.
@@ 6 Conclusion
In this work, we have described a fundamental limitation of many existing embedding initialization methods, which only exploit knowledge stored in a model’s embedding matrices, whereas much of the knowledge about token compositions actually resides in the Transformer layers (Elhage et al., 2022; Lad et al., 2024). To address this limitation, we have proposed Token Distillation, which systematically incorporates this knowledge via distilling contextual information from hidden states. Experimental results confirm that our method not only clearly outperforms “simple” aggregation of embedding matrix rows but also yields better performance than training embeddings using traditional causal language modeling as well as hyper-networks extensively pretrained for new token embedding prediction (Minixhofer et al., 2024). In future work, we believe research into embedding initialization should move beyond an aggregation of information stored in the embedding tables. In particular, a more localized identification of subtoken contextualization (e.g., specific attention heads that aggregate subtokens) is a promising direction for a more targeted distillation objective.
Limitations
@@
Short subword tokens. We find that our method does not work well for lexically ambiguous tokens, i.e., tokens with many different meanings in different contexts. These are often short subwords like <ed> or <fer> that occur in many words. This also means that our method is not ideal to initialize embeddings for an entirely new vocabulary that contains such new subwords. However, these tokens could be initialized using other methods (e.g.Minixhofer et al., 2024), while our method is applied to the rest.
@@
Acknowledgements
We thank the German Federal Ministry for Education and Research (BMBF) for their compute grant through the project <<KI-Servicezentrum Berlin Brandenburg>> (01IS22092). Konstantin Dobler further thanks the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. The research was supported by a research grant (VIL53122) from VILLUM FONDEN, and by the European Union’s Horizon 2020 research and innovation program under grant agreement No. 101135671 (TrustLLM).
Appendix A Additional Results
A.1 Results for “Tokens to Words” (Kaplan et al., 2025)
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Original tokenization 64.5 69.8 70.6 69.2 72.3 60.1 64.4 61.3 66.5 Random (Hewitt, 2021) 57.0±0.6 58.8±0.5 60.6±0.3 59.3±0.5 62.6±0.6 51.6±0.4 55.8±0.2 53.9±0.6 57.5 NTP (tune all embeddings) 55.7±0.5 63.8±0.1 63.8±0.2 62.9±0.4 66.8±0.4 51.5±0.6 55.5±0.5 58.0±0.3 59.8 Mean (Gee et al., 2022) 58.3 63.8 63.4 63.7 66.9 54.2 58.4 58.0 60.8 Tokens to Words (Kaplan et al., 2025) (61.7) (64.3) (65.0) (63.9) (66.6) (57.3) (61.5) (59.5) (62.5) NTP (Lampinen et al., 2018) 61.2±0.4 65.6±0.3 65.3±0.6 66.0±0.5 68.9±0.2 56.6±0.6 61.3±0.5 58.7±0.6 63.0 ZeTT (Minixhofer et al., 2024) 62.7 66.1 66.3 – – – – – – Token Distillation + NTP 63.0±0.5 67.2±0.3 66.7±1.6 67.2±0.3 70.6±0.4 57.6±0.3 62.2±0.3 59.8±0.5 64.3 Token Distillation 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6 Token Distillation + NTP 62.8±0.5 67.6±0.4 67.8±0.5 67.4±0.4 70.9±0.3 57.9±0.1 62.5±0.6 61.2±0.2 64.7
Mistral-7B Llama3-8B Llama3-8B-i Qwen3-8B Avg. Original tokenization 69.5 69.4 72.1 81.7 73.2 Random (Hewitt, 2021) 51.0 46.1 51.5 62.1 52.7 Mean (Gee et al., 2022) 56.3 58.4 61.7 69.6 61.5 Tokens to Words (Kaplan et al., 2025) (64.8) (62.6) (66.6) (82.2) (69.1) NTP (Lampinen et al., 2018) 64.7 67.0 70.1 81.2 70.8 ZeTT (Minixhofer et al., 2024) 66.1 61.3† 65.2† – – Token Distillation 68.5 68.9 72.9 81.5 72.9
@@ Appendix A Additional Results
A.2 Different Target Layers for Token Distillation
@@
Appendix B Implementation Details
B.1 Token Selection for Domain Adaptation
@@
Additionally, to exclude very rare words that would not be commonly used as new tokens, we apply the following two filters: (1) We exclude words that appear fewer than five times across all benchmarks, and (2) we exclude words that occur fewer than 25 times in a sample of 768,000 documents from a domain-specific reference corpus (we use ncbi/pubmed for biomedical benchmarks). The full list of new words is available in our GitHub repository. For illustrative purposes, we provide a random sample of 100 added tokens here:
B.2 Token Selection for French Language Adaptation
We use a similar approach as described in Section B.1 but use HuggingFaceFW/fineweb-2 as a domain-specific reference corpus for filtering rare words and add words that appear in our chosen benchmarks from FrenchBench (Faysse et al., 2025). We also provide a random sample of 100 added tokens for illustrative purposes here:
B.3 Multi-Word Token Generation
@@
B.4 Open Medical-LLM Leaderboard Experiments
@@
This evaluates the task of the Open Medical-LLM Leaderboard (Pal et al., 2024; Jin et al., 2019; 2020; Hendrycks et al., 2021; Pal et al., 2022). All code and a lock-file with specific versions will be available in our GitHub repository.
B.5 FrenchBench Experiments
As for the Open Medical-LLM Leaderboard Experiments, we use the lm-evaluation-harness (Biderman et al., 2024) library at the version 0.4.7 and use 5-shot evaluation on a single Nvidia H100 GPU with bfloat16 precision. We evaluate on the multiple-choice tasks from FrenchBench (Faysse et al., 2025) and use letters (e.g., A, B, C or D) as the answer choices. Specifically, we use the following command:
B.6 LLM-as-a-Judge Experiments
For generating definitions for a particular new token, we use the prompt “The word <new_token> is defined as” following Teehan et al. (2024), where <new_token> is replaced by the string representation of the new token (with whitespace stripped from the left side). We use greedy decoding. Our evaluation considers the model checkpoint corresponding to the best learning rate from the experiments described in Section B.4 for each method.
LLM-as-a-Judge evaluations are not perfect – see for example in Table 2 the similarity scores (Sim) of the target model (measuring similarity with the target model), which should obviously always be 100%, but turn out to be 99.4% / 99.8% in the case of OLMo2-7B-i and Mistral-7B, respectively. Nevertheless, these results (which we report as a sanity check on the judge quality) are very close to correct, which is encouraging.
@@ B.6 LLM-as-a-Judge Experiments
For similarity with the original target model’s definition, we use this prompt:
B.7 Generation of Relevant Contexts
@@
B.8 Isolated Per-Token Optimization vs. Joint Optimization
@@
For the experiments in this paper, we utilize the joint optimization approach. However, we also implemented the isolated optimization of each new embedding for initial experiments. In this setting, gradients are only propagated to a single input embedding (corresponding to the target token of the retrieved snippet), even if other new target tokens also occur in that snippet.
In our exploratory experiments, we found that the joint approach yields slightly better results and conjecture that this is because we are able to have more total gradient updates per target token due to co-occurrence in a particular snippet. As an added benefit, the joint approach has less implementation complexity.
B.9 Hyperparameters

### v3

@@ 5.3 In-Depth Analysis
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Original tokenization 93.4 88.1 99.2 91.1 97.8 65.2 90.9 94.0 90.0 Random (Hewitt, 2021) 0.0 0.0 0.0 0.0 0.0 0.0 0.2 0.0 0.0 Mean (Gee et al., 2022) 2.2 4.2 9.1 8.0 11.3 4.2 7.4 7.4 6.7 NTP Lampinen et al., 2018 34.0 54.3 70.6 49.7 71.6 29.0 53.9 70.0 54.1 ZeTT (Minixhofer et al., 2024) 15.3 25.4 28.4 – – – – – – Token Distillation 46.3 68.0 86.3 61.8 81.3 31.6 57.5 84.3 64.6 Token Distillation + NTP 48.3 68.0 85.3 60.2 79.7 38.2 63.4 80.9 65.5 Token Distillation + NTP 46.9 71.0 85.7 62.0 80.1 37.2 66.4 83.7 66.6
@@
In Table 6, we report LLM-as-a-Judge results on the correctness of generated definitions. This tests the limits of our method, as the new tokens are now more complicated, such as “Software Development” or even common phrases such as “spill the beans”. In this setting, even using the original tokenization sometimes is not able to generate correct definitions. In turn, the gap of Token Distillation to the original tokenization is also larger in this experiment. However, Token Distillation outperforms all other baselines by a very large margin. ZeTT (Minixhofer et al., 2024) in particular struggles in the multi-word token setup, likely because it is out-of-distribution from the hyper-network pretraining.
@@
We find that the subtoken mean initialization as starting point yields better results than random initialization (see “Random + Token Distillation” in Table 8). We use the subtoken mean for its simplicity. In settings where new tokens have low lexical overlap with existing tokens, better results could be achieved via mapping-based initialization methods (e.g., Minixhofer et al., 2022; Dobler & de Melo, 2023). If a pretrained embedding prediction hyper-network is available, we can instead use these predicted embeddings as the starting point. We run this experiment using ZeTT; this combination further improves performance compared to using the subtoken mean (see Table 8). We also compare refining the hyper-network embeddings with NTP. This greatly improves over NTP but yields only minimal gains over using the hyper-network embeddings themselves without further NTP.
@@
Our main proposed method Token Distillation computes the distillation objective based on a mean-squared error (MSE) of last layer hidden states. A natural question is whether we can also use the “traditional” knowledge distillation objective (Hinton et al., 2015) based on the Kullback-Leibler divergence (KL) (Kullback & Leibler, 1951). We analyze this and report the results in Table 9. We additionally include computing the MSE between logits instead of hidden states. For the KL-based results, we include different variations of combining the objective with a NTP-based objective. In general, the choice of distillation objective does not matter much. This is hardly surprising, as the logits are merely a linear projection (potentially with an additional normalization layer) of the hidden state; the log-probabilities used for KL are then in turn merely an additional log-normalization via log-softmax. Combining the KL-based results with NTP-based objectives follows similar trends as their combinations with hidden state MSE-based objectives. We opt for the MSE on hidden states for two reasons: (1) In terms of implementation convenience, the objectives with a projection to logits over the vocabulary require a masking step for the “student” model, as the original “teacher” model does not have the new tokens in its vocabulary. (2) Choosing hidden states allows us to explore other layers than just the last one with potentially large computational speedups – we explore this next.
@@
In our main experiments, we apply the Token Distillation objective on the hidden states after the last layer. However, the last hidden state might not be the most optimal, as it might be overspecialized for next-token prediction (Rogers et al., 2020). In Figure 2, we plot the average performance on our biomedical benchmark datasets using Llama3-8B-Instruct while varying the target layer for extracting hidden states for our Token Distillation objective. Indeed we do see a slight downward trend as we approach the last layer – choosing a different layer than the last one can further boost our already strong results. We repeat the analysis with Llama3.1-8B-Instruct in Section B.3 with matching results.
@@ 6 Conclusion
In this work, we have described a fundamental limitation of many existing embedding initialization methods, which only exploit knowledge stored in a model’s embedding matrices, whereas much of the knowledge about token compositions actually resides in the Transformer layers (Elhage et al., 2022; Lad et al., 2024). To address this limitation, we have proposed Token Distillation, which systematically incorporates this knowledge via distilling contextual information from hidden states. Experimental results confirm that our method not only clearly outperforms methods aggregating embedding matrix rows but also yields better performance than training embeddings using traditional causal language modeling as well as hyper-networks extensively pretrained for new token embedding prediction (Minixhofer et al., 2024). In future work, we believe research into embedding initialization should move beyond an aggregation of information stored in the embedding tables. In particular, a more localized identification of subtoken contextualization (e.g., specific attention heads that aggregate subtokens) is a promising direction for a more targeted distillation objective.
Acknowledgements
We thank the German Federal Ministry of Research, Technology and Space for their compute grant through the project <<KI-Servicezentrum Berlin Brandenburg>> (16IS22092). Konstantin Dobler further thanks the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. The research was supported by a research grant (VIL53122) from VILLUM FONDEN, and by the European Union’s Horizon 2020 research and innovation program under grant agreement No. 101135671 (TrustLLM).
Appendix A Limitations
@@
Short subword tokens. We find that our method does not work as well for lexically ambiguous tokens, i.e., tokens with many different meanings in different contexts. These are often short subwords like <ed> or <fer> that occur in many words. This also means that our method is not ideal to initialize embeddings for an entirely new vocabulary that contains such new subwords. However, these tokens could be initialized using other methods (e.g.Minixhofer et al., 2024), while our method is applied to the rest.
@@
Appendix B Additional Results
B.1 Different Distillation Objectives
We report results for using different distillation objective variants in Table 9. The difference between hidden state or logit-based distillation as well as the differences between MSE and KL based distillation is not significant. However, our chosen approach of hidden state distillation using MSE has the additional advantage of allowing the use of earlier hidden states, offering significant speedups (see Figure 2 and Figure 3).
B.2 Results for “Tokens to Words” (Kaplan et al., 2025)
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Original tokenization 64.5 69.8 70.6 69.2 72.3 60.1 64.4 61.3 66.5 Random (Hewitt, 2021) 57.0±0.6 58.8±0.5 60.6±0.3 59.3±0.5 62.6±0.6 51.6±0.4 55.8±0.2 53.9±0.6 57.5 NTP (tune all embeddings) 55.7±0.5 63.8±0.1 63.8±0.2 62.9±0.4 66.8±0.4 51.5±0.6 55.5±0.5 58.0±0.3 59.8 Mean (Gee et al., 2022) 58.3 63.8 63.4 63.7 66.9 54.2 58.4 58.0 60.8 Tokens to Words (Kaplan et al., 2025) (61.7) (64.3) (65.0) (63.9) (66.6) (57.3) (61.5) (59.5) (62.5) NTP Lampinen et al., 2018 61.2±0.4 65.6±0.3 65.3±0.6 66.0±0.5 68.9±0.2 56.6±0.6 61.3±0.5 58.7±0.6 63.0 ZeTT (Minixhofer et al., 2024) 62.7 66.1 66.3 – – – – – – Token Distillation + NTP 63.0±0.5 67.2±0.3 66.7±1.6 67.2±0.3 70.6±0.4 57.6±0.3 62.2±0.3 59.8±0.5 64.3 Token Distillation 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6 Token Distillation + NTP 62.8±0.5 67.6±0.4 67.8±0.5 67.4±0.4 70.9±0.3 57.9±0.1 62.5±0.6 61.2±0.2 64.7
Mistral-7B Llama3-8B Llama3-8B-i Qwen3-8B Avg. Original tokenization 69.5 69.4 72.1 81.7 73.2 Random (Hewitt, 2021) 51.0 46.1 51.5 62.1 52.7 Mean (Gee et al., 2022) 56.3 58.4 61.7 69.6 61.5 Tokens to Words (Kaplan et al., 2025) (64.8) (62.6) (66.6) (82.2) (69.1) NTP Lampinen et al., 2018 64.7 67.0 70.1 81.2 70.8 ZeTT (Minixhofer et al., 2024) 66.1 61.3† 65.2† – – Token Distillation 68.5 68.9 72.9 81.5 72.9
@@ B.1 Different Distillation Objectives
B.3 Different Target Layers for Token Distillation
@@
B.4 Wall-Clock Speedups from Reduced Over-Tokenization
A note on downstream computational speedups.
Over-tokenization is not only associated with reduced downstream task performance (Rust et al., 2021; Ali et al., 2024) but also with increased computational cost due to longer sequence lengths (Ahia et al., 2023; Yamaguchi et al., 2024a). This computational cost is reduced when reducing over-tokenization by adding new (commonly occurring) tokens to the vocabulary. Importantly, for input embeddings, the reduction in computational cost is completely independent of the actual quality of the embeddings for the newly added tokens (for a given fixed sequence of text). The same computational speedups are achieved by using Random initialization and much better methods such as ZeTT (Minixhofer et al., 2024) or our proposed Token Distillation. Of course, the actual quality of any predictions made conditioned on a sequence containing such new tokens does heavily depend on the embedding quality. Therefore, in this work, we primarily focus on direct evaluation of the quality of produced embeddings. However, we further provide an empirical wall-clock validation of the efficiency gains next.
Empirical speedups.
We empirically validate that reduced token counts translate into concrete wall-clock speedups. We compare batch processing of documents from the arb_Arab split of HuggingFaceFW/fineweb-2 between (i) the original Mistral-7B model and tokenizer and (ii) an adapted model using Token Distillation with 1,000 Arabic tokens selected via AdaptiVocab (Nakash et al., 2025). We repeat the benchmark on both Nvidia H100 GPUs and Nvidia B200 GPUs. We use a single GPU, process 8192 documents per run, and truncate documents to 4k characters. For B200 GPUs, we use a batch size of 4 and for H100 GPUs, we use a batch size of 1 to avoid OOM. We report both inference-style forward-only runtimes and training-style forward+backward runtimes. Note how the compute savings can exceed the token-count reduction due to the quadratic attention cost in sequence length.
Accelerator tokens compute (fwd) compute (fwd+bwd) tok/sec (fwd) tok/sec (fwd+bwd) H100 -35% -41.7% -40.8% +9.5% +7.9% B200 -35% -49.7% -50.5% +24.6% +26.2%
Appendix C Implementation Details
C.1 Token Selection for Domain Adaptation
@@
Additionally, to exclude very rare words that would not be commonly used as new tokens, we apply the following two filters: (1) We exclude words that appear fewer than five times across all benchmarks, and (2) we exclude words that occur fewer than 25 times in a sample of 768,000 documents from a domain-specific reference corpus (we use ncbi/pubmed for biomedical benchmarks). This amounts to roughly 2,600 new tokens (varying per model). The full list of new words is available in our GitHub repository. For illustrative purposes, we provide a random sample of 100 added tokens here:
C.2 Token Selection for French Language Adaptation (Table 3)
We use a similar approach as described in Section C.1 but use HuggingFaceFW/fineweb-2 as a domain-specific reference corpus for filtering rare words and add words that appear in our chosen benchmarks from FrenchBench (Faysse et al., 2025). This amounts to roughly 11,600 new tokens (varying per model). We also provide a random sample of 100 added tokens for illustrative purposes here:
C.3 Multi-Word Token Generation
@@
C.4 Open Medical-LLM Leaderboard Experiments
@@
This evaluates the task of the Open Medical-LLM Leaderboard (Pal et al., 2024; Jin et al., 2019; 2020; Hendrycks et al., 2021; Pal et al., 2022). All code and a lock-file with specific versions are available in our GitHub repository.
C.5 FrenchBench Experiments
As for the Open Medical-LLM Leaderboard Experiments, we use the lm-evaluation-harness (Biderman et al., 2024) library at the version 0.4.7 and use 5-shot evaluation on a single Nvidia H100 GPU with bfloat16 precision. We evaluate on the multiple-choice tasks from FrenchBench (Faysse et al., 2025) and use letters (e.g., A, B, C or D) as the answer choices. We publish the task configuration in our GitHub repository. Specifically, we use the following command:
C.6 Arabic Experiments
We use the lm-evaluation-harness (Biderman et al., 2024) library at the version 0.4.7 and use 5-shot evaluation on a single Nvidia H100 GPU with bfloat16 precision. We evaluate on multiple-choice tasks from the arabic_leaderboard_light (Elfilali et al., 2024) and use letters (e.g., A, B, C or D) as the answer choices. We publish the task configuration in our GitHub repository. Specifically, we use the following command:
C.7 LLM-as-a-Judge Experiments
For generating definitions for a particular new token, we use the prompt “The word <new_token> is defined as” following Teehan et al. (2024), where <new_token> is replaced by the string representation of the new token (with whitespace stripped from the left side). We use greedy decoding. Our evaluation considers the model checkpoint corresponding to the best learning rate from the experiments described in Section C.4 for each method.
LLM-as-a-Judge evaluations are not perfect – see for example in Table 2 the similarity scores (Sim) of the target model (measuring similarity with the original tokenization), which should obviously always be 100%, but turn out to be 99.4% / 99.8% in the case of OLMo2-7B-i and Mistral-7B, respectively. Nevertheless, these results (which we report as a sanity check on the judge quality) are very close to correct, which is encouraging.
@@ C.7 LLM-as-a-Judge Experiments
For similarity with the original model’s definition, we use this prompt:
C.8 Generation of Relevant Contexts
@@
C.9 Isolated Per-Token Optimization vs. Joint Optimization
@@
For the main experiments in this paper, we utilize the joint optimization approach. However, we also implemented the isolated optimization of each new embedding for initial experiments. In this setting, gradients are only propagated to a single input embedding (corresponding to the target token of the retrieved snippet), even if other new target tokens also occur in that snippet.
In our exploratory experiments, we found that the joint approach yields slightly better results and conjecture that this is because we are able to have more total gradient updates per target token due to co-occurrence in a particular snippet. As an added benefit, the joint approach has less implementation complexity. However, when most tokens are actually new tokens (such as in our experiment with adaptation of Mistral-7B-v0.1 to Arabic in Section 5.2), the separate approach should be preferred because the student model mostly sees “noisy” new embeddings instead of just a single one per sequence.
C.10 Hyperparameters